Circular No. 7280

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 T2 F. Shelly reports the discovery by LINEAR of a comet. Additional ob-

servations have been reported following posting on the NEO Confirmation Page:

1999 UT α_{2000} δ_{2000} m_{1} Observer
Oct. 14.15995 22^{h}58^{m}44.^{s}30 +29^{o}22′01′′.4 17.7 LINEAR
14.17563 22 58 43.11 +29 21 56.7 17.7  _{”}
14.19123 22 58 41.85 +29 21 52.8 17.7  _{”}
14.22238 22 58 39.35 +29 21 45.2 18.2  _{”}
14.87973 22 57 47.85 +29 19 00.9  Tichá
14.88149 22 57 47.68 +29 19 00.4  _{”}
14.88353 22 57 47.53 +29 18 59.7 16.5  _{”}
14.88503 22 57 47.41 +29 18 59.7  _{”}
14.91549 22 57 45.04 +29 18 50.9  Busch

LINEAR (Lincoln Laboratory ETS, Socorro). Observers M. Bezpalko, M. Blythe, F. Shelly, M. Elowitz, and R. Huber. Measurers J. Stuart, H. Viggh, and R. Sayer. 1.0-m f/2.15 reflector + CCD.

J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector + CCD. Tail > 20′′

long in p.a. 98^{o}. M. Busch, S. Kluegl, and E. Schwab (Heppenheim). 0.45-m f/4.4 reflector

+ CCD. Tail 10′′ long in p.a. 120^{o}.

SUPERNOVA 1999eg IN IC 1861 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999eg (cf. IAUC 7275), taken by P. Berlind on Oct. 13.4 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova near maximum light. The spectrum exhibits a blue continuum with P-Cyg Balmer lines. Cross-correlation of the host galaxy spectrum with an absorption-line template yields a recession velocity of 6570 ± 50 km/s for IC 1861. The supernova expansion velocity, measured using the Hβ line profile, is 7200 km/s.

COMET C/1997 BA_{6} (SPACEWATCH) Visual m_{1} estimates by A. Pearce, Nedlands, W. Australia (0.41-m

reflector): Sept. 4.88 UT, 12.8; 18.85, 12.8; Oct. 1.86, 12.9; 10.85, 12.7.

1999 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green